### v1

# Doctor-R1: Mastering Clinical Inquiry with Experiential Agentic Reinforcement Learning

Abstract
The professionalism of a human doctor in outpatient service depends on two core abilities: the ability to make accurate medical decisions and the medical consultation skill to conduct strategic, empathetic patient inquiry. Existing Large Language Models (LLMs) have achieved remarkable accuracy on medical decision-making benchmarks. However, they often lack the ability to conduct the strategic and empathetic consultation, which is essential for real-world clinical scenarios. To address this gap, we propose Doctor-R1, an AI doctor agent trained to master both of the capabilities by ask high-yield questions and conduct strategic multi-turn inquiry to guide decision-making. Our framework introduces three key components: a multi-agent interactive environment, a two-tiered reward architecture that separately optimizes clinical decision-making and communicative inquiry skills, and an experience repository to ground policy learning in high-quality prior trajectories. We evaluate Doctor-R1 on OpenAI’s HealthBench and MAQuE, assessed across multi-facet metrics, such as communication quality, user experience, and task accuracy. Remarkably, Doctor-R1 surpasses state-of-the-art open-source specialized LLMs by a substantial margin with higher parameter efficiency and outperforms powerful proprietary models. Furthermore, the human evaluations show a strong preference for Doctor-R1 to generate human-preferred clinical dialogue, demonstrating the effectiveness of the framework. For more details refer to https://github.com/thu-unicorn/Doctor-R1.
Figure 1: The Gap Between Static Medical Desion Task and Dynamic Clinical Inquiry: Top-tiers models demonstrate high performance on static medical decision benchmarks. However, these models adhere to a generic script in a high-risk clinical case, failing to adapt to the complex scenario. Doctor-R1 demonstrates a strategic inquiry process, showing the effectiveness of our framework.
1 Introduction
The outpatient proficiency of a human doctor fundamentally depends on two core abilities: accurate medical decision-making and strategic, empathetic patient inquiry 10.1001/jama.287.2.226. The application of Large Language Models (LLMs) in the medical field has seen impressive advancements in recent years. Frontier and specialized models like GPT-4o (openai2024gpt4ocard) and Med-PaLM 2 (singhal2025expert) have achieved scores surpassing human experts on static medical benchmark like the United States Medical Licensing Examination (USMLE). Building on this foundation, a new wave of advanced medical reasoning models (chen2024huatuogpto1medicalcomplexreasoning; OpenBioLLMs; zhang2024ultramedical; acikgoz2024hippocratesopensourceframeworkadvancing; liu2022mynoserunningarecoughing; xu-etal-2024-reasoning) and multi-agent clinical simulations (li2025agenthospitalsimulacrumhospital; feng2025doctoragentrlmultiagentcollaborativereinforcement; fan-etal-2025-ai) have emerged to tackle more specific clinical challenges.
However, strong performance on static medical benchmarks (jin2020diseasedoespatienthave; jin2019pubmedqadatasetbiomedicalresearch; hendrycks2021measuringmassivemultitasklanguage) does not reflect effective clinical practice. Model performance drops significantly when they are faced with open-ended medical scenarios (kim2025limitationslargelanguagemodels; guan2025keepingmedicalaihealthy; wang2023performancedeteriorationdeeplearning). Real-world clinical practice is not static decision-making task but instead dynamic process of information gathering under uncertainty (baniharouni2025languageagentshypothesisdrivenclinical; helou2020uncertainty). This process which we term Dynamic Inquiry, is a sequential decision-making paradigm where an expert human doctor forms differential diagnostic hypotheses, gathers evidence through targeted questioning to progressively narrow the possibilities, and adapts their inquiry strategy based on the patient’s real-time responses (wilkinson2024oxford), which existing models fail to accomplish (Figure 1).
To assess this capability, a doctor agent should adhere to three key principles: 1) Strategic and Dynamic Inquiry: An effective doctor agent must strategically ask high-yield questions to quickly identify critical risks, rather than following a generic script. For example, in Figure 1, models such as GPT-4.1, HuatuoGPT-o1-70B (chen2024huatuogpto1medicalcomplexreasoning), and DoctorAgent-RL (feng2025doctoragentrlmultiagentcollaborativereinforcement), fails this principle by rigidly following a generic questioning script, rather than adapting its inquiry based on the patient’s responses, resulting in low efficiency and an unsafe consultation. 2) Empathetic Communication: A competent doctor agent must go beyond clinical data collection, communicate with empathy and build patient trust, especially when conveying serious conditions. For instance, Baichuan-M2 (baichuan2025m2) fails this principle when handling emergencies in Table LABEL:tab:inquiry_example. 3) Learning from Good Experience: An advanced doctor agent should learn from prior experiences like a real physician, by selectively retrieving high-quality experiences to continuously refine its inquiry strategy. Existing models fail to support dynamic inquiry (Principle 1), empathic communication (Principle 2), and high-quality experience learning (Principle 3), which relying primarily on similarity-based retrieval (yan2025memoryr1enhancinglargelanguage; zhong2023memorybankenhancinglargelanguage; li2025agenthospitalsimulacrumhospital).
To this end, we propose Doctor-R1, a novel framework that adheres to these three principles. In particular, 1) To enable dynamic inquiry, Doctor-R1 leverages a Reinforcement Learning (RL) framework with Group Relative Policy Optimization (GRPO) (shao2024deepseekmathpushinglimitsmathematical) in a multi-agent interactive environment for multi-turn dialogue training. 2) To integrate empathetic communication, we introduced a two-tiered reward architecture which consists of process reward and outcome reward. The reward function is defined through multi-dimensional metrics that evaluate both hard and soft skills of the agent. The policy is directly optimized by the process reward that explicitly scores for empathy. 3) To learn from good experience, we introduce an Experiential Agentic Reinforcement Learning (RL) approach to guide the agent’s decisions. We construct an experience repository that selectively stores and retrieves high-quality prior trajectories based on their rewards. A multi-stage experience retrieval pipeline is used to ensure the agent learns from the most relevant, novel, and high-reward experiences, which we define as the characteristics of a good experience.
We evaluate Doctor-R1 on two challenging medical benchmarks, HealthBench (arora2025healthbenchevaluatinglargelanguage) and MAQuE (MAQuE), covering a wide range of medical specialties and complex multi-turn diagnostic scenarios. The results demonstrate that Doctor-R1 outperforms all other open-source specialized LLMs, establishing a new state-of-the-art for doctor agents. Doctor-R1 outperforms Baichuan-M2 by a significant margin on HealthBench communication scores +25.05 (47.16 vs. 22.11) by conducting more efficient questioning. Specifically, Doctor-R1 observed a substantial improvement in with an increase of +9.91 (36.29. vs. 26.38) on HealthBench, and +8.00 (60.00 vs. 52.00) on MAQuE compared to UltraMedical-70B (chen2024huatuogpto1medicalcomplexreasoning). Notably, our 8B model also significantly surpasses powerful proprietary LLMs like GPT-4.1, Grok 4, Claude Sonnet 4, and Gemini 2.5 Flash, demonstrating that our Experiential Agentic RL framework is more effective than model scale for mastering the dynamic clinical consultations.
The main contributions of this work are as follows:
  • We introduce Doctor-R1, the first agent framework to unify two core clinical skills: strategic multi-turn inquiry (soft skills) and medical decision-making (hard skills) within a single agent.
  • We propose a new closed-loop Experiential Agentic Reinforcement Learning methodology which integrates a multi-agent interactive environment, a two-tiered reward architecture that separately optimizes decision-making and communicative skills, and an experience repository guiding policy learning from high-quality prior trajectories.
  • Doctor-R1 surpasses top-tier open-source and proprietary LLMs on HealthBench and MAQuE with greater parameter efficiency, highlighting a key finding: enhancing inquiry capability improves decision-making. These results are further validated by strong human preferences in evaluation studies.
2 Preliminaries
Agentic Reinforcement Learning (RL) (zhang2025landscapeagenticreinforcementlearning) refers to a paradigm focused on training LLMs as autonomous goal-directed agents rather than static conditional generators optimized for single-turn alignment or benchmark performance. Unlike traditional supervised fine-tuning which focuses on imitation, Agentic RL situates the policy model ($\pi_{\theta}$) in a partially observable, interactive environment where it learns sequences of actions toward long-term objectives. Guided by a reward model ($\mathcal{R}_{\psi}$) and updated via optimization algorithms, Agentic RL is exceptionally well-suited for complex sequential tasks like medical inquiry, as it trains the agent to develop a strategic policy for actively gathering information and making decisions under uncertainty.
Policy Model
In the context of Agentic RL, the policy model is the agent being trained to generate an action $a_{t}$ (an utterance) given an observation $o_{t}$ (the dialogue history). The policy, denoted as $\pi_{\theta}$ with parameters $\theta$, is typically a large language model whose objective is to learn a strategy that maximizes rewards from the environment. In this work, the doctor agent is the policy model.
Multi-Objective Reward Model
A reward model $\mathcal{R}_{\psi}$ provides a proxy signal for evaluating policy actions. In complex tasks like medical consultations, a single reward cannot capture diverse goals such as accuracy, efficiency, and empathy. We adopt a multi-objective reward model, a vector-valued function $\mathcal{R}_{\psi}(o_{t},a_{t})\in\mathbb{R}^{K}$, where $o_{t}$ is the observation, $a_{t}$ the action, $\psi$ the parameters, and $K$ the number of reward dimensions which evaluate different aspects of quality. In our framework, the Consultation Evaluator functions as the multi-objective reward model.
GRPO Model
GRPO (Group Relative Policy Optimization) algorithm (shao2024deepseekmathpushinglimitsmathematical) optimizes the policy model $\pi_{\theta}$ by leveraging the relative quality of multiple candidate responses within a group. Instead of relying on a single advantage estimate like PPO or a single chosen/rejected pair like DPO, GRPO uses a listwise objective based on reward scores across a response set. Specifically, GRPO optimizes the policy to prefer a chosen response $y_{c}$ over a group of rejected responses $\{y_{r,j}\}_{j=1}^{N}$, by maximizing of the log-likelihood of the chosen response relative to the group:
$\mathcal{L}_{\text{GRPO}}=-\mathbb{E}_{(x,y_{c},\{y_{r}\})\sim\mathcal{D}}\left[\log\frac{e^{R_{\psi}(x,y_{c})}}{\sum_{y\in\{y_{c}\}\cup\{y_{r}\}}e^{R_{\psi}(x,y)}}\right]$ (1)
where $R_{\psi}(x,y)$ is the reward score. This loss function encourages higher reward for $y_{c}$ relative to all rejections. In our framework, the policy model $\pi_{\theta}$ is the doctor agent, and the multi-objective reward model $R_{\psi}$ is the Consultation Evaluator. We apply GRPO to optimize the the policy of the Doctor Agent by contrasting a high-quality action against a diverse group of less effective actions.
3 Doctor-R1
In this section, we describe the construction of Doctor-R1, a framework designed for continuous experiential reinforcement learning for doctor agents. We begin by defining the dynamic interactive environment in Section 3.1, which specifies how the training environment is constructed with Markov decision process in agents. Next, in Section 3.2 we present our two-tiered reward architecture which evaluates both conversational process (soft skills) and final diagnostic correctness (hard skills). Finally, we detail the experience repository that enables the agent to store and retrieve high-quality prior trajectories to ground its policy in Section 3.3.
Figure 2: The interactive training loop of our Doctor-R1 framework. The process unfolds within a (1) Dynamic Interactive Environment populated by diverse patient simulations. At (2) each turn of an inquiry session, the (3) Doctor Agent interacts with the environment by observing the state, queries the (4) Experience Repository, and selects an action. A Patient Agent responds, and the (5) Consultation Evaluator evaluates the action based on the two-tiered reward architecture. This new experience is stored into the repository, and is used to optimize the policy of Doctor Agent.
3.1 Dynamic Interactive Environment
Following the Agentic RL paradigm, we formalize the medical consultation process as a Partially Observable Markov Decision Process (POMDP). This framework is well-suited to model the sequential, uncertain nature of clinical inquiry, where a doctor agent cannot directly observe the true state (i.e., the full medical condition of patient) but must act based on a history of observations. To construct this environment, we define three core interacting components that form a closed-loop feedback system. The environment’s dynamics are governed by the POMDP tuple $\langle\mathcal{S,A,O,R}\rangle$, with the following components: 1) States ($\mathcal{S}$): A state $s\in\mathcal{S}$ represents the consultation query of the patient agent. 2) Actions ($\mathcal{A}$): An action $a_{t}\in\mathcal{A}$ is an utterance generated by the doctor agent at turn $t$. 3) Observations ($\mathcal{O}$): An observation $o_{t}\in\mathcal{O}$ is the dialogue history up to turn $t$. At each step, the doctor agent observes the patient’s latest response and adds it to the history to form the new observation $o_{t+1}$. 4) Reward Function ($\mathcal{R}$): Reward $\mathcal{R}(s,a)$ provides a signal to the doctor agent after each action. It is implemented by the Consultation Evaluator which evaluates action $a_{t}$ as detailed in Section 3.2. Each element is realized through the interaction of these components:
  • Doctor Agent ($\pi_{\theta}$): This is our target agent. Its policy, $\pi_{\theta}(a_{t}|o_{t})$, maps the current observation $o_{t}$ (dialogue history) to a distribution over actions $a_{t}\in\mathcal{A}$. The goal is to learn an optimal inquiry policy $\pi_{\theta}^{*}$ that maximizes the cumulative reward.
  • Patient Agent: This agent simulates patient behavior, effectively defining the environment’s state transitions. Portrayed by an independent LLM, it is initialized with a clinical scenario that defines its underlying true state $s\in S$. It generates a response based on the action $a_{t}$, which updates the next observation $o_{t+1}$ for the Doctor Agent.
  • Consultation Evaluator: This component acts as the environment’s reward function $\mathcal{R}$. At each turn $t$, it provides a multi-dimensional reward by assessing the action in the dialogue context.
This closed-loop feedback system provides the foundation for training the polify of the doctor agent. Doctor agent is prompted to learn how to ask “good questions” in the multi-turn conversation loop until it finally comes to a diagnostic conclusion, or reaches the predefined maximum conversation turns. The LLM-based patient agent acts as a core component in the interactive environment, providing diverse responses and dynamically constructing a high diversity environment enriched with various clinical tasks, enable the doctor agent to interact, learn, and evolve through each rollout.
3.2 Two-Tiered Reward Architecture
To enable the robust learning of the agent over long-horizons, a simple monolithic reward is insufficient (gao2022scalinglawsrewardmodel). We introduce a two-tiered reward architecture that evaluates the two core clinical skills of a physician: both the conversational process (soft skills) and the final diagnostic outcome (hard skills). The process reward gives turn-by-turn feedback to shape the agent behaviour, while the outcome reward evaluates the global accuracy of the entire dialogue.
Process Reward for Conversational Quality
To shape the inquiry capabilities of the agent, we provide a dense process reward ($R_{\text{turn}}$) after each turn. This reward evaluates the quality of the conversation, focusing on communication, safety, and reasoning. The agent responses are evaluated across eight dimensions, $S_{i}$: 1) Safety: Penalizes harmful advice or unsafe recommendation. 2) Reasoning: The logical and medical soundness of the reasoning process. 3) Medical Accuracy: The factual correctness of information presented to the patient. 4) Completeness: Fully addresses user concerns and provides clear, useful next steps. 5) Information Gathering: The effectiveness of the questions in assessing urgency and relevance. 6) Faithfulness to Ground Truth: Alignment with expert-provided standards for the given scenario. 7) Empathy and Clarity: The use of simple and reassuring language. 8) Humility: Appropriate language and penalizes unwarranted certainty.
Our design is built on a “safety-first” principle, using a hierarchical penalty to handle critical failures. This addresses the limitations of conventional weighted-sum models, which can fail to adequately penalize catastrophic errors due to an “averaging-out” effect (amodei2016concreteproblemsaisafety). As shown in Equation 2, any violation of safety, reasoning, or accuracy triggers an immediate large negative reward, overriding the standard score calculation. This veto system conceptually aligned with shielded frameworks in Safe RL, where a safety layer intervenes to prevent unsafe actions (Alshiekh_Bloem_Ehlers_Könighofer_Niekum_Topcu_2018), ensuring that fundamental clinical reliability is non-negotiable. If no violation is triggered, the reward is a weighted sum of all eight dimensional scores.
$R_{\text{turn}}=\begin{cases}-1.0&\text{if }S_{\text{safety}}<\epsilon\\ -0.75&\text{else if }S_{\text{reasoning}}<\epsilon\text{ or }S_{\text{accuracy}}<\epsilon\\ \sum_{i=1}^{k}w_{i}\cdot S_{i}&\text{otherwise}\end{cases}$ (2)
Here, $\epsilon$ is the failure threshold, $S_{i}$ is the score for each of the eight dimensions, and $W_{i}$ are the corresponding weights that reflect their relative importance. The specific weights and normalization details are further elaborated in Appendix LABEL:appendix:reward.
Outcome Reward for Diagnostic Accuracy
At the end of a dialogue episode, the agent receives a single terminal outcome reward ($R_{\text{final}}$) based on the correctness of the final diagnosis. This reward evaluates the core “hard skill” of the agent, which is the ability of decision-making throughout the consultation. The decision of the agent is compared against a ground-truth diagnosis and assigns a score ($S_{c}orrectness$) based on the degree of alignment. This ensures the agent is optimized not just to interact well, but to drive the conversation towards a medically sound conclusion. A score of 1.0 is given for a correct diagnosis, 0.5 for a partially correct one (e.g., correct differential but wrong primary diagnosis), and 0.0 for an incorrect diagnosis. For both reward types, the model is required to generate a reasoning trace before scoring to ensure more consistent judgments.
$R_{\text{final}}=S_{\text{correctness}}\quad\text{where}\quad S_{\text{correctness}}\in\{0.0,0.5,1.0\}$ (3)
3.3 Experience Repository
Multi-stage Experience Retrieval
The doctor agent optimizes its policy from past trajectories (state, action, reward) using a multi-stage retrieval pipeline, where we combine an embedding model for candidate selection and a reranker model for fine-grained reordering, balancing the trade-off between low-precision embedding and computationally expensive reranking.
  • Stage I. Candidate Selection: The objective of this stage is to efficiently reduce the search space of experiences. The query state $Q$ and all stored experience states $E_{i}^{\text{state}}$ are precomputed into dense embeddings. The combined score $S_{\text{combined}}$ (Equation 4) integrates cosine similarity with the historical reward $R(E_{i})$ weighted by $\alpha$. The top-$\mathcal{N}$ entries are selected to form $E_{\text{candidates}}\subset E$, narrowing the database from millions of experiences to a manageable number for the next stage. Here, $f_{\text{emb}}$ is the embedding function and $f_{\text{sim}}$ is the cosine similarity function.
    $S_{\text{combined}}(Q,E_{i})=f_{\text{sim}}(f_{\text{emb}}(Q),f_{\text{emb}}(E_{i}^{\text{state}}))+\alpha\cdot R(E_{i})$ (4)
  • Stage II. High-Fidelity Reranking: The candidate set $E_{\text{candidates}}$ is then passed to a more powerful reranker model. While the bi-encoder architecture of the embedding model processes the query independently, the reranker model uses a cross-encoder architecture to score each pair $(Q,E_{i}^{\text{state}})$ with token-level attention for more accurate relevance, and reorder the candidates accordingly.
  • Stage III. Novelty and Reward Filtering: The novelty filtering mechanism prevents the agent from repeatedly retrieving highly similar experiences, and the reward filtering mechanism retains high-performing experiences. For the reranked candidate set, a dynamic high-reward threshold $\tau_{\text{dynamic}}$ is calculated as in Equation 5, where $\mu_{R}$ and $\sigma_{R}$ are the mean and standard deviation of rewards within the candidate set, and $\beta_{\text{std}}$ is a configurable factor. This prunes the candidate experience set based on the combination of the semantic similarity to the query and the reward.
    $\tau_{\text{dynamic}}=\mu_{R}(E_{\text{candidates}})+\beta_{\text{std}}\cdot\sigma_{R}(E_{\text{candidates}})$ (5)
    $E_{j}\in E\Leftrightarrow\left(f_{\text{sim}}(f_{\text{emb}}(Q),f_{\text{emb}}(E_{i}^{\text{state}}))<\tau_{\text{novelty}}\right)\quad\land\quad\left(R(E_{i})>\tau_{\text{dynamic}}\right)$ (6)
    An experience $E_{j}\in E_{\text{candidates}}$ is considered as a “good experience”, and added to the full experience set $E$, if two conditions are met: (1) its semantic similarity to the query $Q$ falls below a predefined novelty threshold $\tau_{\text{novelty}}$, and (2) its associated reward $R(E_{i})$ exceeds the dynamic high-reward threshold (Equation 6). The retrieved top-$k$ experiences are then prepended to the original query for the doctor agent to provide a direct suggestion for the next move of the agent.
Selective Experience Storage
To enable the agent learning from valuable experiences, high-reward interactions are selectively stored as experience tuples $E_{t}=(s_{t},a_{t},R_{t})$, where $s_{t}$ is the state, $a_{t}$ the action, and $R_{t}$ the reward. Following the processing of each batch $\mathcal{B}$, an experience $E_{i}$ is stored if and only if its reward $R_{i}$ meets or exceeds a predefined high-reward boundary, $\tau_{\text{reward}}$. The filtered set of high-reward experiences, denoted as $E_{\text{new}}=\{E_{i}\in\mathcal{B}\mid R_{i}\geq\tau_{\text{reward}}\}$. This ensures the long-term preservation of valuable knowledge across training sessions. The in-memory cache and embedding tensor are updated in real time for immediate retrieval and long-term retention.
4 Evaluation Results
Our primary evaluations are conducted on HealthBench (arora2025healthbenchevaluatinglargelanguage) and MAQuE (MAQuE), two benchmarks featuring complex clincal scenarios and multi-turn dialogues. To validate that Doctor-R1’s specialized training does not compromise its foundational knowledge, we also evaluate it on the static QA benchmark MedQA (jin2020diseasedoespatienthave) and MMLU (hendrycks2021measuringmassivemultitasklanguage), to confirm the robustness of our model’s knowledge base.
4.1 Overall Performance
Table LABEL:tab:healthbench_main shows the overall performance on HealthBench Main, which is evaluated across two dimensions: 1) Themes: categorize the medical scenario (Emergency Referrals, Health Data Tasks, Communication, Global Health, Hedging, Context Seeking, and Complex Responses), and 2) Axes: measure fundamental skills (Accuracy, Communication Quality, Instruction Following, Context Awareness, and Completeness). The detailed result of HealthBench Hard is listed in Appendix LABEL:appendix:detailed_results.
Table 1: Overall model performance on HealthBench Main (best results of open-source models are bolded)